One year​ ago, your company purchased a machine used in manufacturing for

$ 105 comma 000$105,000.

You have learned that a new machine is available that offers many​ advantages; you can purchase it for

$ 160 comma 000$160,000

today. It will be depreciated on a​ straight-line basis over ten​ years, after which it has no salvage value. You expect that the new machine will contribute EBITDA​ (earnings before​ interest, taxes,​ depreciation, and​ amortization) of

$ 35 comma 000$35,000

per year for the next ten years. The current machine is expected to produce EBITDA of

$ 25 comma 000$25,000

per year. The current machine is being depreciated on a​ straight-line basis over a useful life of 11​ years, after which it will have no salvage​ value, so depreciation expense for the current machine is

$ 9 comma 545$9,545

per year. All other expenses of the two machines are identical. The market value today of the current machine is

$ 50 comma 000$50,000.

Your​ company's tax rate is

40 %40%​,

and the opportunity cost of capital for this type of equipment is

12 %12%.

Is it profitable to replace the​ year-old machine?

The NPV of the replacement is

​$nothing.

​(Round to the nearest​ dollar.)

Please show me how you got the answer
Answer #1

New Machine Depreciation = 160,000/10 = 16,000

Old Machine Depreciation = 105,000/11 = 9,545

Year 0 Year 1-10
Gross Margin    10,000.00
Depreciation      6,455.00
EBIT      3,545.00
Tax      1,418.00
Incremental Earnings      2,127.00
Add:Depreciation      6,455.00
Capex - 160,000.00
     68,182.00
FCF -   91,818.00      8,582.00

NPV = -91,818 + (8,582 x 5.6502)

= -43,327.98

No its not profitable to replace the old machine